OpenStudy (anonymous):

simplify the expression 10P6/10P4

OpenStudy (anonymous):

p^2

OpenStudy (anonymous):

I dont have any options like that my options are 420 151200 30 5040

OpenStudy (anonymous):

@cmckie0208

OpenStudy (anonymous):

oh sorry I thought the 6 and 4 were exponents

OpenStudy (anonymous):

no haha. do you know how to do this?

OpenStudy (anonymous):

@Hero can you help?

hero (hero):

\(nPr = \dfrac{n!}{(n - r)!}\)

OpenStudy (anonymous):

so... 10/(10-4)??

OpenStudy (anonymous):

@Hero

hero (hero):

\[10P6 = \frac{10!}{(10 - 6)!}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

wait so I got 2.5?

hero (hero):

Did you find \(\dfrac{10P6}{10P4}\) ?

OpenStudy (anonymous):

I got 2.5 idk what i did wrong

hero (hero):

Do you know what n factorial is and how to evaluate it?

OpenStudy (anonymous):

no I dont

hero (hero):

\(n! = n(n - 1)(n - 2)(n - 3)...(3)(2)(1)\) For example \(6! = (6)(5)(4)(3)(2)(1) = 720\)

hero (hero):

Do you understand that?

OpenStudy (anonymous):

To be honest not really lol. Could you show me step by step to solve this expression? I learn so much better when I can see each step if you dont mind

hero (hero):

Okay, I'll show you an example \(\dfrac{7P5}{7P3} = \dfrac{\dfrac{7!}{(7 - 5)!}}{\dfrac{7!}{(7 - 3)!}}\) \(= \dfrac{\dfrac{7!}{2!}}{\dfrac{7!}{4!}} \\= \dfrac{\dfrac{(7)(6)(5)(4)(3)(2)(1)}{(2)(1)}}{\dfrac{(7)(6)(5)(4)(3)(2)(1)}{(4)(3)(2)(1)}} \\= \dfrac{\dfrac{(7)(6)(5)(4)(3)\cancel{(2)(1)}}{\cancel{(2)(1)}}}{\dfrac{(7)(6)(5)\cancel{(4)(3)(2)(1)}}{\cancel{(4)(3)(2)(1)}}} \\ = \dfrac{(7)(6)(5)(4)(3)}{(7)(6)(5)} \\ = \dfrac{\cancel{(7)(6)(5)}(4)(3)}{\cancel{(7)(6)(5)}} \\ = (4)(3) \\ = 12\)

OpenStudy (anonymous):

Thank you! The answer I got was 30... is that correct?

hero (hero):

Correct

OpenStudy (anonymous):

thank you!